OpenStudy (anonymous):

Evaluate. 3 – 7 • 5 + 2 • 4 A. –40 B. –24 C. –12 D. 28

OpenStudy (anonymous):

@KamiBug

OpenStudy (kamibug):

B. -24

OpenStudy (kamibug):

3 – 7 • 5 + 2 • 4 3 – 35 + 8 -32 + 8 -24

OpenStudy (anonymous):

thanks \[\frac{ 3.2+1.4\div7 }{9.3-7.6\times4 }\]

OpenStudy (kamibug):

\[\frac{ 3.4 }{ -21.1 }\]

OpenStudy (anonymous):

the bottom is wrong it will not go in @KamiBug

OpenStudy (kamibug):

What do you mean? The negative sign? It can also be in the numerator, it doesn't change the value. :)

OpenStudy (anonymous):

ok can you please hold on one second

OpenStudy (kamibug):

Sure

OpenStudy (anonymous):

Evaluate the expression, using the given value of the variable. 3 – 7x + 2x when x = 3 A. –12 B. –24 C. –30 D. –18

OpenStudy (kamibug):

3 – 7x + 2x when x = 3 Just plug in the 3. 3 - 7(3) + 2(3) 3 - 21 + 6 -24

OpenStudy (anonymous):

Evaluate the expression, using the given value of the variable. z ÷ 3 – 5z + 4(z – 2) when z = 12 A. –16 B. –14 C. –96 D. –98

OpenStudy (anonymous):

@KamiBug

OpenStudy (kamibug):

z ÷ 3 – 5z + 4(z – 2) when z = 12 Same thing. You just gotta use substitution. :) What does the equation look like when you plug in the value of z?

OpenStudy (anonymous):

ok is it B?

OpenStudy (anonymous):

I do not know the answer i think -14 though?

OpenStudy (anonymous):

@KamiBug

OpenStudy (kamibug):

It's C. -96

OpenStudy (anonymous):

hold on please

OpenStudy (anonymous):

it was -16 right?

OpenStudy (kamibug):

It's C. -96

OpenStudy (anonymous):

kk one sec i have one more:)

OpenStudy (kamibug):

Okay

OpenStudy (anonymous):

What is \[\left(\begin{matrix}n-5\times6 \\ n \div4\end{matrix}\right)\]when n = 8?

OpenStudy (kamibug):

Is it a fraction?

OpenStudy (anonymous):

i can not get this plllllllleeeeeeeaaaaaaaaasssssssssssseeeeeeeeeeee help

OpenStudy (anonymous):

no

OpenStudy (anonymous):

it is but the answer is not a fraction:)

OpenStudy (kamibug):

Oh, yeah that's what I meant. :P Gime a sec.

OpenStudy (anonymous):

i meaN K:)

OpenStudy (kamibug):

\[\frac{ 8-5*6 }{ 8/4 }\]

OpenStudy (anonymous):

YES

OpenStudy (kamibug):

What's 8 minus 5 times 6?

OpenStudy (anonymous):

22

OpenStudy (anonymous):

5*6 is 30-8=22

OpenStudy (kamibug):

Well, remember the 8 is in the front. So you're doing 8-30 not 30-8. So it's guna be -22 :) Now, what's 8 divided by 4?

OpenStudy (anonymous):

2

OpenStudy (kamibug):

Right! :) So in the numerator we have -22, and the denominator is 2. So our fraction is now -22/2. And we can divide :) What is -22 / 2 ?

OpenStudy (anonymous):

-11:)

OpenStudy (kamibug):

You've got it!:D

OpenStudy (anonymous):

i got a 20% i failed it is ok but thanks:)

OpenStudy (kamibug):

How many questions were there in total?